Can you help me solve this Linear Equation
x-5y=-154
x-3y=-84

Answer by WannabeCAgirl83(35)


There you go... ;-)

x – 5y = - 154 (Lets name this Equation 1)
x – 3y = - 84 (Lets name this Equation 2)

Step # 1
Isolating x in Equation 1:

x – 5y = - 154
x – 5y = - 154 | + 5y on both sides
x = - 154 + 5y

Step # 2
Substituting x in Equation 2 with the term found for x in step # 1, then isolating y :

x – 3y = - 84
- 154 + 5y – 3y = - 84 | re-arranging
- 154 + 2y = - 84 | + 154 on both sides
2y = 70 | : 2 on both sides
y = 35

Step # 3
Substituting y in Equation 1 with the term found for y in step # 2, then again isolating x :

x – 5y = - 154
x – 5(35) = - 154
x – 175 = - 154 | + 175 on both sides
x = 21

Step # 4:
Validating solutions by substituting them with x and y in both Equations:

x – 5y = - 154
21 – 5(35) = -154 → true

x – 3y = - 84
21 – 3(35) = - 84 → true